Quality control is an important issue at ACME Company, which manufactures light bulbs. To test the life-hours of their light bulbs, they randomly sampled nine light bulbs and measured how many hours they lasted: 378, 361, 350, 375, 200, 391, 375, 368, 321.
What is the mean?
Answer #1

Mean = (Sum of observations) / (Total number of observations)

Here,

Sum of observations = 378 + 361 + 350 + 375 + 200 + 391 + 375 + 368 + 321 = 3119

Total number of observations = 9

Therefore,

Mean = 3119/9 = 346.56